A companys balance sheet contalned the following Information: $13,5ee Total Assets Common stock Accounts Payable $191,000 29,50e 65,500 Retained Earnings Notes Payable Is the only other item on the balance sheet Notes Payable must equal: Multiple Cnoice $82.500. $17,000. 584.400 $19,300.
Answer #1

((191,000)= L (Account payable 65,500+ Notes payable x)+ SE (Retained Earnings 29,500+ Common Stock 13,500)
191,000 = 65,500 + x+29,500 + 13,500
191,000 = 108,500 + x
191,000-108,500=x
82,500=x)

So The Answer is $82,500